# Hierarchical Clustering With Confidence

### v1

@@ 6 Conclusion
There are several promising directions for future work. In this paper, we have derived exact p-values under normality. Extending our approach to distribution-free settings, future work could investigate the construction of asymptotically valid p-values by using the transfer of conditional validity to other distributions or non-parametric settings, of which has been established in recent work on randomized decision trees (Bakshi and Panigrahi, 2025). Moreover, although our current approach focuses on hierarchical clustering algorithms, the potential of our randomization scheme to develop inference for other types of clustering methodology as well as for other types of hypotheses about differences between clusters remains to be explored.
Appendix A Extensions
A.1 Constructing p-value under non-spherical covariance
@@
For the observed realization of the data matrix $X=X_{o}$ and an observed merge $\{M^{(t)}=M_{o}^{(t)}\}$, where $M_{o}^{(t)}=\left(C^{(t)}_{1,o},C^{(t)}_{2,o}\right)$, we consider the null hypothesis in Gao et al. [2024], defined as:
$H_{0}^{(t)}:\bar{\mu}_{C^{(t)}_{1,o}}=\bar{\mu}_{C^{(t)}_{2,o}}$ (19)
where $\bar{\mu}_{C}=\frac{1}{|C|}\sum_{i\in C}\mu_{i}$ is defined as the mean for cluster $C$.
A.1.1 Test statistics and preliminaries
@@ A.1.2 Computing the p-value
$\mathbb{F}^{(t)}(u;\mathcal{A}^{(t)}_{\Sigma,o},\overline{M}_{o}^{(t)})=\frac{\int_{0}^{u}\ell_{\|\nu_{M_{o}^{(t)}}\|_{2}\cdot\chi_{p}}(u^{\prime})\times\prod_{s=1}^{t}p^{(s)}\left(M_{o}^{(s)};X(u^{\prime};\mathcal{A}_{\Sigma,o}^{(t)})\right)du^{\prime}}{\int_{0}^{\infty}\ell_{\|\nu_{M_{o}^{(t)}}\|_{2}\cdot\chi_{p}}(u^{\prime})\times\prod_{s=1}^{t}p^{(s)}\left(M_{o}^{(s)};X(u^{\prime};\mathcal{A}_{\Sigma,o}^{(t)})\right)du^{\prime}},$
where $\ell_{\|\nu_{M_{o}^{(t)}}\|_{2}\cdot\chi_{p}}(u)$ denotes the density of a $\|\nu_{M_{o}^{(t)}}\|_{2}\cdot\chi_{p}$ random variable and $p^{(s)}\left(M_{o}^{(s)};X(u^{\prime};\mathcal{A}_{\Sigma,o}^{(t)})\right)$ is the probability that $M^{(s)}=M_{o}^{(s)}$ given $X=X(u^{\prime};\mathcal{A}_{o}^{(t)})$, for $s\in\{1,2,\ldots,t\}$, as defined in (1).
@@
Furthermore, since the merges were sampled according to the probabilities in (1), and the sampling processes were independent given the data, we have that
@@
$=\prod_{s=1}^{t}\mathbb{P}\left(M^{(s)}=M_{o}^{(s)}|X=X(u;\mathcal{A}_{\Sigma,o}^{(t)})\right)=\prod_{s=1}^{t}p^{(s)}\left(M_{o}^{(s)};X(u;\mathcal{A}_{\Sigma,o}^{(t)})\right).$
@@
$\ell_{\|\nu_{M_{o}^{(t)}}\|_{2}\cdot\chi_{p}}(u)\times\prod_{s=1}^{t}p^{(s)}\left(M_{o}^{(s)};X(u;\mathcal{A}_{\Sigma,o}^{(t)})\right).$
@@
$\frac{\ell_{\|\nu_{M_{o}^{(t)}}\|_{2}\cdot\chi_{p}}(u)\times\prod_{s=1}^{t}p^{(s)}\left(M_{o}^{(s)};X(u;\mathcal{A}_{\Sigma,o}^{(t)})\right)}{\int_{0}^{\infty}\ell_{\|\nu_{M_{o}^{(t)}}\|_{2}\cdot\chi_{p}}(u^{\prime})\times\prod_{s=1}^{t}p^{(s)}\left(M_{o}^{(s)};X(u^{\prime};\mathcal{A}_{\Sigma,o}^{(t)})\right)du^{\prime}}.$
@@
${\textbf{P}}^{(t)}_{\Sigma,o}=\mathbb{P}_{H_{0}^{(t)}}\left(U^{(t)}\geq U^{(t)}_{o}\big\lvert\left\{\overline{M}^{(t)}=\overline{M}_{o}^{(t)},\mathcal{A}^{(t)}_{\Sigma}=\mathcal{A}^{(t)}_{\Sigma,o}\right\}\right)$,

### v2

@@ 6 Conclusion
There are several promising directions for future work. In this paper, we have derived exact p-values under normality. To extend our approach to distribution-free settings, future work could investigate the construction of asymptotically valid p-values by using the transfer of conditional validity to other distributions or to non-parametric settings, similar to what has recently been done for randomized decision trees [Bakshi and Panigrahi, 2025]. Moreover, although our current approach focuses on hierarchical clustering algorithms, the potential of our randomization scheme to develop inference for other types of clustering methodology and for other types of cluster-based hypotheses remains to be explored.
APPENDIX
[…]
We first present a preliminary result establishing that, at a given step $t$, the random merge made by our algorithm converges to the set of admissible merges in traditional clustering, conditional on the previous history of merges and the observed data.
Lemma 3.
Suppose $(\overline{M}^{*(t-1)}_{o},M^{*(t)}_{o})\in\Omega_{o}^{*(t)}$ be a merge sequence from traditional hierarchical clustering applied to a fixed data matrix $X_{o}$. Then, we have that
$\lim_{\tau^{*}\rightarrow 0}\mathbb{P}((\overline{M}^{*(t-1)}_{o},M^{(t)})\in\Omega_{o}^{*(t)}|\overline{M}^{(t-1)}=\overline{M}^{*(t-1)}_{o},X=X_{o})=1.$
Proof.
By definition of traditional hierarchical clustering,
$d(M^{*(t)}_{o};X_{o})<d(M;X_{o}),\quad\forall\quad M\notin\underset{M^{\prime}\in\mathcal{M}(\overline{M}^{*(t-1)}_{o})}{\text{argmin}}d(M^{\prime},X_{o}).$
Now, observe the following:
$\mathbb{P}(M^{(t)}=M|\overline{M}^{(t-1)}=\overline{M}^{*(t-1)}_{o},X=X_{o})$ $=\dfrac{\exp\left(-\dfrac{d(M;X_{o})}{\tau(\mathcal{M}^{(t)}_{o},X_{o})}\right)}{\displaystyle\sum_{M^{\prime}\in\mathcal{M}^{(t)}_{o}}\exp\left(-\dfrac{d(M^{\prime};X_{o})}{\tau(\mathcal{M}^{(t)}_{o},X_{o})}\right)}\times\dfrac{\exp\left(\dfrac{d(M^{*(t)}_{o};X_{o})}{\tau(\mathcal{M}^{(t)}_{o},X_{o})}\right)}{\exp\left(\dfrac{d(M^{*(t)}_{o};X_{o})}{\tau(\mathcal{M}^{(t)}_{o},X_{o})}\right)}$
$=\dfrac{\exp\left(-\dfrac{1}{\tau(\mathcal{M}^{(t)}_{o},X_{o})}(d(M;X_{o})-d(M^{*(t)}_{o};X_{o}))\right)}{\displaystyle\sum_{M^{\prime}\in\mathcal{M}^{(t)}_{o}}\exp\left(-\dfrac{1}{\tau(\mathcal{M}^{(t)}_{o},X_{o})}(d(M^{\prime};X_{o})-d(M^{*(t)}_{o};X_{o}))\right)}$
$=\frac{\exp\left(-\dfrac{|\mathcal{M}^{(t)}_{o}|}{\tau^{*}}\times\dfrac{(d(M;X_{o})-d(M^{*(t)}_{o};X_{o}))}{\sum_{M^{\prime}\in\mathcal{M}^{(t)}_{o}}d(M^{\prime};X_{o})}\right)}{1+\displaystyle\sum_{M^{\prime}\in\mathcal{M}^{(t)}_{o}\setminus M^{*(t)}_{o}}\exp\left(-\dfrac{|\mathcal{M}^{(t)}_{o}|}{\tau^{*}}\times\dfrac{(d(M;X_{o})-d(M^{*(t)}_{o};X_{o}))}{\sum_{M^{\prime}\in\mathcal{M}^{(t)}_{o}}d(M^{\prime};X_{o})}\right)}$
$<\exp\left(-\dfrac{|\mathcal{M}^{(t)}_{o}|}{\tau^{*}}\times\dfrac{(d(M;X_{o})-d(M^{*(t)}_{o};X_{o}))}{\sum_{M^{\prime}\in\mathcal{M}^{(t)}_{o}}d(M^{\prime};X_{o})}\right).$
As $\tau^{*}\rightarrow 0$, the right-hand side of the above inequality tends to 0 since $d(M;X_{o})-d(M^{*(t)}_{o};X_{o})>0$. Hence,
$\mathbb{P}(M^{(t)}=M|\overline{M}^{(t-1)}=\overline{M}^{*(t-1)}_{o},X=X_{o})\rightarrow 0$
for all $M\notin\underset{M^{\prime}\in\mathcal{M}(\overline{M}^{*(t-1)}_{o})}{\text{argmin}}d(M^{\prime},X_{o})$. Since the number of possible merges is finite, it follows that
$\lim_{\tau^{*}\to 0}\mathbb{P}(M^{(t)}\in\underset{M^{\prime}\in\mathcal{M}(\overline{M}^{*(t-1)}_{o})}{\text{argmin}}d(M^{\prime},X_{o})\mid\overline{M}^{(t-1)}=\overline{M}^{*(t-1)}_{o},X=X_{o})=1,$
which is equivalent to
$\lim_{\tau^{*}\rightarrow 0}\mathbb{P}((\overline{M}^{*(t-1)}_{o},M^{(t)})\in\Omega_{o}^{*(t)}|\overline{M}^{(t-1)}=\overline{M}^{*(t-1)}_{o},X=X_{o})=1.$
Now, we are ready to present to proof for Proposition 1.
Proof.
We prove this result by induction. When $t=1$, we have that
$\mathbb{P}(\overline{M}^{(1)}\in\Omega^{*(1)}_{o}|X=X_{o})\rightarrow 1$
as $\tau^{*}\rightarrow 0$, directly from Lemma 3.
Now suppose that the claim holds up to step t1t-1, i.e.
$\lim_{\tau^{*}\rightarrow 0}\mathbb{P}(\overline{M}^{(t-1)}\in\Omega^{*(t-1)}_{o}|X=X_{o})=1,$ (17)
At step $t$, by the law of total probability,
$\quad\mathbb{P}(\overline{M}^{(t)}\in\Omega^{*(t)}_{o}|X=X_{o})$
$=\sum_{\overline{M}^{*(t-1)}_{o}\in\Omega_{o}^{*(t-1)}}\mathbb{P}((\overline{M}^{*(t-1)}_{o},M^{(t)})\in\Omega_{o}^{*(t)},\ \overline{M}^{(t-1)}=\overline{M}^{*(t-1)}_{o}\mid X=X_{o})$
$=\sum_{\overline{M}^{*(t-1)}_{o}\in\Omega^{*(t-1)}_{o}}\mathbb{P}((\overline{M}^{*(t-1)}_{o},M^{(t)})\in\Omega^{*(t)}_{o}\mid\overline{M}^{(t-1)}=\overline{M}^{*(t-1)}_{o},X=X_{o})\times\mathbb{P}(\overline{M}^{(t-1)}=\overline{M}^{*(t-1)}_{o}\mid X=X_{o})$
By Lemma 3, for each fixed $\overline{M}^{*(t-1)}_{o}\in\Omega^{*(t-1)}_{o}$,
$\lim_{\tau^{*}\rightarrow 0}\mathbb{P}((\overline{M}^{*(t-1)}_{o},M^{(t)})\in\Omega_{o}^{*(t)}\mid\overline{M}^{(t-1)}=\overline{M}^{*(t-1)}_{o},\ X=X_{o})=1.$
Therefore, it holds that
$\lim_{\tau^{*}\rightarrow 0}\mathbb{P}(\overline{M}^{(t)}\in\Omega^{*(t)}_{o}|X=X_{o})$ $=\lim_{\tau^{*}\rightarrow 0}\sum_{\overline{M}^{*(t-1)}_{o}\in\Omega^{*(t-1)}_{o}}\mathbb{P}(\overline{M}^{(t-1)}=\overline{M}^{*(t-1)}_{o}\mid X=X_{o})$
$=\lim_{\tau^{*}\rightarrow 0}\mathbb{P}(\overline{M}^{(t-1)}\in\Omega^{*(t-1)}_{o}\mid X=X_{o})=1,$
where the second equality comes from the definition of $\Omega^{*(t-1)}_{o}$ in (4) and the last equality comes from (17). This completes the proof. ∎
[…]
Proof.
Note that the observed data matrix $X_{o}$ can be decomposed as follows:
$X_{o}$ $=B^{(t)}_{o}X_{o}+W^{(t)}_{o}X_{o}+(I_{n}-B^{(t)}_{o}-W^{(t)}_{o})X_{o}$
The right-hand side of the above display further equals
[…]
Proof.
Moreover, under the null hypothesis, we have $\mathbb{E}[B^{(t)}_{o}X]=\mathbb{E}[W^{(t)}_{o}X]=0$. By the properties of the Gaussian distribution with zero mean and isotropic covariance, it follows that $(\|B^{(t)}_{o}X\|_{F},\|W^{(t)}_{o}X\|_{F})\perp\!\!\!\perp\left(\eta^{(t)},\gamma^{(t)}\right)$. Therefore, $R^{(t)}$ as a statistic that is a measurable function of $\|B^{(t)}_{o}X\|_{F}$, $\|W^{(t)}_{o}X\|_{F}$ remains independent of $\left(\eta^{(t)},\gamma^{(t)}\right)$.
$R^{(t)}\Big\lvert\mathcal{A}^{(t)}=\mathcal{A}_{o}^{(t)}$
[…]
We first present a preliminary result that derives the probability of the observed merge sequence conditional on the test statistic and the auxiliary statistics.
Lemma 4.
For $t\geq 1$, we have that
$\mathbb{P}\!\left(M^{(1)}=M_{o}^{(1)},\ldots,M^{(t)}=M_{o}^{(t)}\mid R^{(t)}=r,\;\mathcal{A}^{(t)}=\mathcal{A}^{(t)}_{o}\right)=\prod_{s=1}^{t}p^{(s)}\left(M_{o}^{(s)};\mathcal{M}_{o}^{(s)},X(r;\mathcal{A}_{o}^{(t)})\right)$
where
$p^{(s)}\left(M_{o}^{(s)};\mathcal{M}_{o}^{(s)},X(r;\mathcal{A}_{o}^{(t)})\right)=\mathbb{P}\left(M^{(s)}=M_{o}^{(s)}|\overline{M}^{(s-1)}=\overline{M}_{o}^{(s-1)},X=X(r;\mathcal{A}_{o}^{(t)})\right)$
is the sampling probability that $M^{(s)}=M_{o}^{(s)}$ given $X=X(r;\mathcal{A}_{o}^{(t)})$ and the merge history $\overline{M}^{(s-1)}=\overline{M}_{o}^{(s-1)}$, for $s\in\{1,2,\ldots,t\}$, as defined in (2).
Proof.
By Lemma 1, we have that
$\mathbb{P}\left(M^{(1)}=M_{o}^{(1)},\ldots,M^{(t)}=M_{o}^{(t)}\;|\;R^{(t)}=r,\mathcal{A}^{(t)}=\mathcal{A}_{o}^{(t)}\right)$ (18)
$=\mathbb{P}\left(M^{(1)}=M_{o}^{(1)},\ldots,M^{(t)}=M_{o}^{(t)}\;\mid X=X(r;\mathcal{A}_{o}^{(t)})\right).$
By the chain rule of probability, one can further write the probability on the right-hand side of the above-stated equality as
$\mathbb{P}(M^{(1)}=M^{(1)}_{o}\mid X=X(r;\mathcal{A}_{o}^{(t)}))\times\prod_{s=2}^{t}\mathbb{P}(M^{(s)}=M_{o}^{(s)}\mid\overline{M}^{(s-1)}=\overline{M}_{o}^{(s-1)},X=X(r;\mathcal{A}_{o}^{(t)}))$
From the definition of our sampling scheme in (2), it follows that the merge variable $M^{(s)}$, conditional on $\overline{M}^{(s-1)}=\overline{M}_{o}^{(s-1)}$ and $X=X(r;\mathcal{A}_{o}^{(t)})$, takes the value $M_{o}^{(s)}$ with probability $p^{(s)}\left(M_{o}^{(s)};\mathcal{M}_{o}^{(s)},X(r;\mathcal{A}_{o}^{(t)})\right)$, i.e.
$\mathbb{P}(M^{(s)}=M_{o}^{(s)}\mid\overline{M}^{(s-1)}=\overline{M}_{o}^{(s-1)},X=X(r;\mathcal{A}_{o}^{(t)}))=p^{(s)}\left(M_{o}^{(s)};\mathcal{M}_{o}^{(s)},X(r;\mathcal{A}_{o}^{(t)})\right).$
Hence, the probability in the claim factorizes as
$\mathbb{P}\!\left(\bigcap_{s=1}^{t}\{M^{(s)}=M_{o}^{(s)}\}\mid X=X(r;\mathcal{A}_{o}^{(t)})\right)=\prod_{s=1}^{t}p^{(s)}\left(M_{o}^{(s)};\mathcal{M}_{o}^{(s)},X(r;\mathcal{A}_{o}^{(t)})\right).$
Now, we are ready to present the proof for Theorem 1.
Proof.
Let the conditional density of $R^{(t)}\Big\lvert\mathcal{A}^{(t)}=\mathcal{A}_{o}^{(t)}$ at $r$ be denoted by $\ell_{R^{(t)}}(r\;\lvert\;\mathcal{A}^{(t)}=\mathcal{A}_{o}^{(t)}).$ Then, using Bayes’ rule, we observe that the density of the conditional distribution $R^{(t)}\Big\lvert\{\overline{M}^{(t)}=\overline{M}_{o}^{(t)},\mathcal{A}^{(t)}=\mathcal{A}_{o}^{(t)}\},$ evaluated at $r$, is proportional to
Using Lemma 2, under the null $H_{0}^{(t)}$, it follows that
Furthermore, by Lemma 4, we have that
$=\mathbb{P}\left(\bigcap_{s=1}^{t}\{M^{(s)}=M_{o}^{(s)}\}\mid X=X(r;\mathcal{A}_{o}^{(t)})\right)=\prod_{s=1}^{t}p^{(s)}\left(M_{o}^{(s)};\mathcal{M}_{o}^{(s)},X(r;\mathcal{A}_{o}^{(t)})\right).$
Combining (20) with (21), we obtain that the expression in (19) takes the value $\ell_{F_{p,(N_{o}^{(t)}-2)p}}(r)\times\prod_{s=1}^{t}p^{(s)}\left(M_{o}^{(s)};\mathcal{M}_{o}^{(s)},X(r;\mathcal{A}_{o}^{(t)})\right).$ That is, the conditional density of interest is given by
$\frac{\ell_{F_{p,(N_{o}^{(t)}-2)p}}(r)\times\prod_{s=1}^{t}p^{(s)}\left(M_{o}^{(s)};\mathcal{M}_{o}^{(s)},X(r;\mathcal{A}_{o}^{(t)})\right)}{\int_{0}^{\infty}\ell_{F_{p,(N_{o}^{(t)}-2)p}}(r^{\prime})\times\prod_{s=1}^{t}p^{(s)}\left(M_{o}^{(s)};\mathcal{M}_{o}^{(s)},X(r^{\prime};\mathcal{A}_{o}^{(t)})\right)dr^{\prime}}.$
[…]
Proof.
Fixing some notation, let
Then, by the probability integral transform, we obtain that
$\mathbb{P}_{\overline{M}_{o}^{(t)}}(\mathcal{A}^{(t)}_{o})=\alpha.$
Finally, we apply the law of iterated expectation to conclude that
[…]
Proof.
By Lemma 3, for each step $1\leq s\leq t$, it holds almost everywhere that, as $\tau^{*}\rightarrow 0$,
$p^{(s)}\left(M_{o}^{(s)};\mathcal{M}_{o}^{(s)},X(r;\mathcal{A}_{o}^{(t)})\right)\rightarrow\mathbbm{1}\{M_{o}^{(s)}\in\Psi^{*(s)}(r)\},$
where $\Psi^{*(s)}(r)=\displaystyle\operatorname*{arg\,min}_{M\in\mathcal{M}^{(s)}_{o}}d(M;X(r;\mathcal{A}^{(s)}_{o}))$.
Hence, the weight function
$w_{\tau^{*}}(r):=\prod_{s=1}^{t}p^{(s)}\left(M_{o}^{(s)};\mathcal{M}_{o}^{(s)},X(r;\mathcal{A}_{o}^{(t)})\right)$
converges almost everywhere to $\prod_{s=1}^{t}\mathbbm{1}\{M_{o}^{(s)}\in\Psi^{*(s)}(r)\}$. Furthermore, the weight functions are uniformly bounded, i.e., $0\leq w_{\tau^{*}}(r)\leq 1$. Since the integrands in the numerators and denominators of $\mathbb{F}^{(t)}(R^{(t)}_{o};\mathcal{A}^{(t)}_{o},\overline{M}_{o}^{(t)})$, derived in Theorem 1, converge almost everywhere, respectively, to the integrands of (15), applying the dominated convergence theorem to the numerator and denominator of $\mathbb{F}^{(t)}(R^{(t)}_{o};\mathcal{A}^{(t)}_{o},\overline{M}_{o}^{(t)})$ yields the stated limit. ∎
[…]
Lemma 5.
$\mathbb{P}(\mathcal{V})\leq\alpha.$
Proof.
For each step $t$, consider the event
Observe that
We are now ready to present the proof of Theorem 3 using Lemma 5.
Theorem 3.
The proof of this result follows from observing that
[…]
@@
For the observed realization of the data matrix $X=X_{o}$ and an observed merge $\{M^{(t)}=M_{o}^{(t)}\}$, where $M_{o}^{(t)}=\left(C^{(t)}_{1,o},C^{(t)}_{2,o}\right)$, we consider the null hypothesis in Gao et al. [2024], defined as
$H_{0}^{(t)}:\bar{\mu}_{C^{(t)}_{1,o}}=\bar{\mu}_{C^{(t)}_{2,o}}$, (23)
where $\bar{\mu}_{C}=\frac{1}{|C|}\sum_{i\in C}\mu_{i}$ is defined as the mean of the population means in cluster $C$.
[…]
@@
$\mathbb{F}^{(t)}(u;\mathcal{A}^{(t)}_{\Sigma,o},\overline{M}_{o}^{(t)})=\frac{\int_{0}^{u}\ell_{\|\nu_{M_{o}^{(t)}}\|_{2}\cdot\chi_{p}}(u^{\prime})\times\prod_{s=1}^{t}p^{(s)}\left(M_{o}^{(s)};\mathcal{M}_{o}^{(s)},X(u^{\prime};\mathcal{A}_{\Sigma,o}^{(t)})\right)du^{\prime}}{\int_{0}^{\infty}\ell_{\|\nu_{M_{o}^{(t)}}\|_{2}\cdot\chi_{p}}(u^{\prime})\times\prod_{s=1}^{t}p^{(s)}\left(M_{o}^{(s)};\mathcal{M}_{o}^{(s)},X(u^{\prime};\mathcal{A}_{\Sigma,o}^{(t)})\right)du^{\prime}},$
where $\ell_{\|\nu_{M_{o}^{(t)}}\|_{2}\cdot\chi_{p}}(u)$ denotes the density of a $\|\nu_{M_{o}^{(t)}}\|_{2}\cdot\chi_{p}$ random variable and $p^{(s)}\left(M_{o}^{(s)};\mathcal{M}_{o}^{(s)},X(u^{\prime};\mathcal{A}_{\Sigma,o}^{(t)})\right)$ is the probability that $M^{(s)}=M_{o}^{(s)}$ given the merge history $\overline{M}^{(s-1)}=\overline{M}_{o}^{(s-1)}$ and $X=X(u^{\prime};\mathcal{A}_{o}^{(t)})$, for $s\in\{1,2,\ldots,t\}$, as defined in (2).
@@
Furthermore, by Lemma 4, we have that
@@
$=\mathbb{P}\left(\bigcap_{s=1}^{t}\{M^{(s)}=M_{o}^{(s)}\}|X=X(u;\mathcal{A}_{\Sigma,o}^{(t)})\right)=\prod_{s=1}^{t}p^{(s)}\left(M_{o}^{(s)};\mathcal{M}_{o}^{(s)},X(u;\mathcal{A}_{\Sigma,o}^{(t)})\right).$
@@
$\ell_{\|\nu_{M_{o}^{(t)}}\|_{2}\cdot\chi_{p}}(u)\times\prod_{s=1}^{t}p^{(s)}\left(M_{o}^{(s)};\mathcal{M}_{o}^{(s)},X(u;\mathcal{A}_{\Sigma,o}^{(t)})\right).$
@@
$\frac{\ell_{\|\nu_{M_{o}^{(t)}}\|_{2}\cdot\chi_{p}}(u)\times\prod_{s=1}^{t}p^{(s)}\left(M_{o}^{(s)};\mathcal{M}_{o}^{(s)},X(u;\mathcal{A}_{\Sigma,o}^{(t)})\right)}{\int_{0}^{\infty}\ell_{\|\nu_{M_{o}^{(t)}}\|_{2}\cdot\chi_{p}}(u^{\prime})\times\prod_{s=1}^{t}p^{(s)}\left(M_{o}^{(s)};\mathcal{M}_{o}^{(s)},X(u^{\prime};\mathcal{A}_{\Sigma,o}^{(t)})\right)du^{\prime}}.$
@@
${\textbf{P}}^{(t)}_{\Sigma,o}=\mathbb{P}_{H_{0}^{(t)}}\left(U^{(t)}\geq U^{(t)}_{o}\big\lvert\overline{M}^{(t)}=\overline{M}_{o}^{(t)},\mathcal{A}^{(t)}_{\Sigma}=\mathcal{A}^{(t)}_{\Sigma,o}\right)$,